# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

### v2

@@ Appendix A Details of Rating Algorithms
Purpose: is used to create a partial order based on the computed weighted rejection score / the PIE %.
@@
$S$, Set of systems.
$P$, Set of perturbations.
$F$, Flag that says whether the confounder is present (1) or not (0).
$D$, $CI$, $W$ (as defined in the previous algorithms).
@@
$PO$, dictionary with a partial order for each perturbation.
$PO\leftarrow\{\}$;
@@
if F == 0 then
       for each $p_{i}\in P$ do
             for each $s_{j}\in S$ do
                   $\psi\leftarrow WeightedRejectionScore(p_{i},s_{j},D,CI,W)$;
                   $SD[s_{j}]\leftarrow\psi$;
             end for
            $PO[p_{i}]\leftarrow SORT(SD)$;
       end for
else
      for each $p_{j}\in P$ do
             for each $s_{i}\in S$ do
                   $\psi\leftarrow ComputePIEScore(s_{i},p_{j},p_{0},D,CI,W)$;
                   $SD[s_{j}]\leftarrow\psi$;
             end for
            $PO[p_{i}]\leftarrow SORT(SD)$;
@@
return $PO$
Algorithm 3 CreatePartialOrder
Purpose: AssignRating assigns a rating to each of the SASs based on the partial order and the number of rating levels, $L$.
Input:
$S$, $D$, $CI$, $W$, $P$ (as defined in the previous algorithms).
$L$, rating levels chosen by the user.
Output:
$R$, dictionary with perturbations as keys and ratings assigned to each system within each perturbation as the values.
$R\leftarrow\{\}$;
$PO\leftarrow\text{CreatePartialOrder}(S,D,CI,W,G)$;
for $p_{i}\in P$ do
[…]
      if $\text{len}(S)>1$ then
             $G\leftarrow\text{ArraySplit}(\psi,L)$;
             for $k,i\in PO[p_{i}],\psi$ do
                   for $g_{j}\in G$ do
                         if $i\in g_{j}$ then
                               $SD[k]\leftarrow j$;
                         end if
                   end for
             end for

### v3

@@ 4.3. Workflow
Purpose: Calculate the Average Treatment Effect (ATE).
@@
$s$, $d$, $p$, $p_{0}$ (as defined in the previous algorithm)
@@
$\psi$, ATE score.
$\psi\leftarrow E(R^{max}_{t}|do(P=p))-E(R^{max}_{t}|do(P=p_{0}))$      // Compute ATE.
return $\psi$
Algorithm 3 ComputeATEScore
Purpose: Create a partial order of systems within each treatment based on raw scores (APE/ATE, PIE, or WRS).
Input:
$S$, $d$ (as defined in the previous algorithm); $P$, Set of treatments; MetricMetric, specifies which metric to use (APE/ATE, PIE, or WRS).
Output:
POPO, dictionary with partial orders for each treatment.
$PO\leftarrow\{\}$      // Initialize dictionary for partial orders.
$SD\leftarrow\{\}$
for each $p_{i}\in P$ do
  $SD\leftarrow\{\}$      // Reset scores for each treatment.
 for each $s_{j}\in S$ do
     if $Metric==\text{APE/ATE}$ then
        $\psi\leftarrow ComputeATEScore(s_{j},d,p_{i},p_{0})$      // Compute ATE.
    else
        if $Metric==\text{PIE}$ then
           $\psi\leftarrow ComputePIEScore(s_{j},d,p_{i},p_{0})$      // Compute PIE.
       else
           $\psi\leftarrow WeightedRejectionScore(p_{i},s_{j},d)$      // Compute WRS.
        end if
     end if
    $SD[s_{j}]\leftarrow\psi$      // Store score for system $s_{j}$.
  end for
 $PO[p_{i}]\leftarrow SORT(SD)$      // Sort systems in ascending order of scores.
end for
return POPO
Algorithm 4 CreatePartialOrder
Purpose: Assign a rating to each system based on the partial order and the number of rating levels, $L$.
Input:
$S$, $d$, $P$, MetricMetric (as defined in the previous algorithm); $L$, rating levels chosen by the user.
Output:
$R$, dictionary with perturbations as keys and ratings for each system as values.
$R\leftarrow\{\}$;
$PO\leftarrow\text{CreatePartialOrder}(S,d,P,Metric)$;
for $p_{i}\in P$ do
  $\psi\leftarrow[PO[p_{i}].\text{values()}]$;
 if $\text{len}(S)>1$ then
     $G\leftarrow\text{ArraySplit}(\psi,L)$;
@@
     for $k,i\in PO[p_{i}]$ do
        $SD[k]\leftarrow\text{FindGroup}(i,G)$;
@@
FindGroup Function:
Input: valuevalue, groupsgroups; Output: Group index.
for $g_{j}\in groups$ do
  if $value\in g_{j}$ then
     return index of $g_{j}$;
  end if
end for